# An Actionable Framework for Assessing Bias and Fairness in Large Language Model Use Cases

### v1

@@ 1 Introduction
The versatility of current Large Language Models (LLMs) in handling various tasks (Minaee et al., 2024; Liu et al., 2023; Ray, 2023) presents challenges when it comes to evaluating bias and fairness at the model level. Existing approaches primarily focus on assessing risk using benchmark data sets containing predefined prompts (Gehman et al., 2020; Dhamala et al., 2021; Nozza et al., 2021; Smith et al., 2022; Parrish et al., 2021; Li et al., 2020), masked tokens (Zhao et al., 2018; Rudinger et al., 2018; Nadeem et al., 2021; Levy et al., 2021), or unmasked sentences (Nangia et al., 2020; Barikeri et al., 2021; Jiao et al., 2023; Felkner et al., 2023), assuming that these adequately capture specific bias or fairness risks (Gallegos et al., 2023). However, these assessments are likely to overestimate the risk for use cases where the population of prompts is low risk (Wang et al., 2024). Moreover, to the best of the author’s knowledge, the current literature does not provide a framework for effectively aligning LLM use cases with suitable metrics for evaluating bias and fairness.
@@
This work aims to address these limitations by developing an actionable LLM bias and fairness evaluation framework defined at the use case level. Drawing inspiration from the classification fairness framework proposed by Saleiro et al. (2018), the framework proposed in this work enables practitioners to map an LLM use case to an appropriate set of bias and fairness evaluation metrics by considering relevant characteristics of the use case and stakeholder values. This evaluation approach is unique in that it incorporates actual prompts from the practitioner’s use case, taking into account the prompt-specific risks that have been demonstrated to significantly increase the likelihood of biased and unfair outcomes (Wang et al., 2024). By constraining the scope to focused use cases, where prompts are derived from a known population and the task is well-defined, this framework is specifically designed to customize the risk assessment for a specific application.
@@
To introduce the framework, this study first provides formal definitions of bias and fairness desiderata for LLMs from the literature and segment these definitions by risk category. Subsequently, those risks are mapped to a taxonomy of use cases focused on large-scale applications, where human-in-the-loop may be infeasible. Lastly, for each risk category, various bias and fairness evaluation metrics are detailed, with discussions provided on their input requirements, calculation methods, the risks they assess, and circumstances under which they should be applied. As part of this work, a variety of novel bias and fairness metrics are introduced. This includes innovative counterfactual adaptations of recall-oriented understudy for gisting evaluation (ROUGE) (Lin, 2004), bilingual evaluation understudy (BLEU) (Papineni et al., 2002), and cosine similarity (Singhal and Google, 2001), as well a set of stereotype classifier-based metrics that are adapted from analogous toxicity classifier-based metrics.
@@ 2 Bias and Fairness Risks for LLM Use Cases
This section formally defines several pre-requisite terms and concepts upon which subsequent sections rely, several of which are adapted from those provided by Gallegos et al. (2023). [^2] The notation and terminology used in this paper differs from those used by Gallegos et al. (2023). The approach outlined below adopts modified versions of these definitions to better fit the tone and purpose. For further reading on biases in LLMs, the reader may refer to (Gallegos et al., 2023; Blodgett et al., 2020; Kumar et al., 2023; Li et al., 2024; Chu et al., 2024; Ferrara, 2023; Ranaldi et al., 2023; Kotek et al., 2023; Wu and Aji, 2023; Li et al., 2023; Ray, 2023; Nozza et al., 2022; Zhuo et al., 2023). The discussions in this paper explore concepts of bias and fairness in relation to an arbitrary ‘protected attribute’, encompassing examples such as sex, race, age, and sexual orientation, among others. These concepts are established formally by using mathematical notation whenever possible and strives to prioritize consistency in notation throughout, even if it may differ from the notation used in the papers cited.
@@ 2.2.3 Counterfactual Fairness
In many contexts, it is undesirable for an LLM to generate substantially different output as a result of different protected attribute words contained in the input prompts, all else equal (Huang et al., 2020; Nozza et al., 2021; Wang et al., 2024). Following previous work (Huang et al., 2020; Garg et al., 2019), this concept is hereafter referred to as (lack of) counterfactual fairness. Depending on context and stakeholder values, the practitioner may wish to assess an LLM for differences in overall content or sentiment resulting from inclusion of different protected attribute words in a prompt. Below, this section provides a formal definition of the corresponding fairness desiderata known as counterfactual invariance, adapted from Gallegos et al. (2023).
@@ 3 Bias and Fairness Evaluation Metrics
Accounting for prompt-specific risk is important to accurately reflect the risk of a specific use case, as Wang et al. (2024) find in their evaluations that toxicity probability is 26 to 101 times higher when dealing with toxic prompts compared to non-toxic prompts. Moreover, Goldfarb-Tarrant et al. (2020); Delobelle et al. (2021) have demonstrated that evaluation metrics which take into account the LLM’s task provide a more accurate reflection of the associated risk compared to metrics based on embeddings or token probabilities. Accordingly, the objective of this work is to assess the risks associated with a particular use case, taking into consideration not only the LLM used but also the task at hand and the prompt population. Following the characterization of a use case based on a model and a known population of prompts, as established in Section 2.1, each metric definition presented below is contextualized within an evaluation sample of size $N$ drawn from a known population of prompts $\mathcal{P}_{X}$.
@@ 3.1 Bias Metrics for Text Generation and Summarization Use Cases
While Gallegos et al. (2023) further segment generated-text metrics into three subcategories (namely distribution metrics, classifier metrics, and lexicon metrics), this work segments these metrics according to the risk taxonomy outlined in 2.2. In particular, generated-text bias and fairness metrics are segmented into the following categories: toxicity metrics, stereotype metrics, and counterfactual fairness metrics. Toxicity metrics leverage a pre-trained toxicity classifier, such as Perspective API [^12] https://perspectiveapi.com, to assign a toxicity score to an LLM’s output (Chowdhery et al., 2022; Lees et al., 2022; Wang et al., 2024; Bommasani et al., 2023; Gehman et al., 2020). Stereotype metrics assess the relative co-occurrence of stereotype words with protected attribute words (Bordia and Bowman, 2019; Bommasani et al., 2023) or leverage a pre-trained stereotype classifier to assign a stereotype score to an LLM’s output (Zekun et al., 2023). Lastly, counterfactual fairness metrics assess fairness through a causal lens by exploiting variation in mentions of protected attributes in otherwise-identical prompts (Huang et al., 2020). Select metrics are detailed in each category below. [^13] Note that the discussion of generated text bias metrics in this paper is not exhaustive. Rather, the intention is to select an appropriate set of evaluation metrics to address each risk to be assessed. Notable metrics omitted from this framework include HONEST (Nozza et al., 2021), score parity (Sicilia and Alikhani, 2023), Full Gen Bias (Smith et al., 2022), regard score (Sheng et al., 2019), demographic representations (Bommasani et al., 2023), psycholinguistic norms (Dhamala et al., 2021), and gender polarity (Dhamala et al., 2021).
@@ 3.1.1 Toxicity Metrics
Toxicity metrics leverage a pre-trained toxicity classifier to assess general toxicity that is not specific to a particular protected attribute. Given a toxicity classifier $T:\mathcal{Y}\xrightarrow[]{}[0,1]$, this section defines toxicity metrics in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$. Specifically, these metrics are evaluated on a set of $m$ responses per prompt, $\{\hat{Y}_{ij}\}_{j=1}^{m}$, generated by $\mathcal{M}$, from a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$. In practice, variation in responses for the same prompt can be achieved by choosing a non-zero temperature for the LLM or by utilizing various random seeds (Wang et al., 2024). Below, formal definitions of three toxicity metrics are provided: Expected Maximum Toxicity, Toxicity Probability, and Toxic Fraction. All three metrics range in values from 0 to 1, with smaller values indicating a greater degree of fairness.
@@
Note that the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024).
@@
As with EMT, the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024).
@@ 3.1.2.1 Co-occurrence-Based Metrics
Given two protected attribute groups $G^{\prime},G^{\prime\prime}$ with associated sets of protected attribute words $A^{\prime},A^{\prime\prime}$, a set of stereotypical words $W$, and an LLM use case $(\mathcal{M},\mathcal{P}_{X})$, the full calculation of COBS is as follows:
$P(w|A)=\frac{\sum_{i=1}^{N}cooccur(w,A|\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\hat{Y}_{i}}cooccur(\tilde{w},A|\tilde{Y}_{i})}{\sum_{i=1}^{N}\sum_{a\in A}C(a,\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\tilde{Y}_{i}}C(\tilde{w},\hat{Y}_{i})}$
$COBS=\frac{1}{|W|}\sum_{w\in W}\log\frac{P(w|A^{\prime})}{P(w|A^{\prime\prime})}.$
Above, $C(x,\hat{Y}_{i})$ denotes the count of $x$ in $\hat{Y}_{i}$ and $\tilde{Y}_{i}$ represents the LLM output $\hat{Y}_{i}$, with words from $A^{\prime}\cup A^{\prime\prime}$ and stop words excluded. The co-occurrence function $cooccur(w,A|\hat{Y})$ computes a weighted count of words from $A$ that are found within a context window centered around $w$, each time $w$ appears in $\hat{Y}$. [^15] To specify the co-occurrence function $cooccur(w,A|\hat{Y})$, Bordia and Bowman (2019) define a fixed or infinite context window centered at a target word $w$. In the case of fixed context (Bordia and Bowman (2019) recommend including 10 words before and 10 words after), each word in the context window receives equal weight. In the case of an infinite context window, each word receives weight $\beta^{dist},$ where $dist$ denotes the number of tokens between the two tokens of interest (Bordia and Bowman (2019) use $\beta=0.95$). In both cases, only words from the context window included in $A$ are included in the weighted count. In words, COBS computes the relative likelihood that an LLM $\mathcal{M}$ generates output having co-occurrence of $w\in W$ with $A^{\prime}$ versus $A^{\prime\prime}$. This metric has a range of possible values of [,][$\infty$,$\infty$][ - ∞ , ∞ ], with values closer to 0 signifying a greater degree of fairness.
@@
Consider a set of protected attribute groups $\mathcal{G}$, an associated set of protected attribute lexicons $\mathcal{A}$, and an associated set of stereotypical words $W$. Additionally, let $C(x,\hat{Y})$ denote the number of times that the word $x$ appears in the output $\hat{Y}$, $I(\cdot)$ denote the indicator function, $P^{\text{ref}}$ denote a reference distribution, and $TVD$ denote total variation difference. [^16] The reference distribution recommended by Bommasani et al. (2023) is the uniform distribution. Total variation distance measures the distance between probability distributions. For a given LLM $\mathcal{M}(X;\theta)$ and a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$, the full computation of SA is as follows:
@@ 3.1.3.1 Counterfactual Similarity
Counterfactual similarity metrics measure the similarity in outputs generated from counterfactual input pairs according to a specified invariance metric $\upsilon$, i.e. $\upsilon(\mathcal{M}(X^{\prime};\theta),\mathcal{M}(X^{\prime\prime};\theta)).$ These metrics effectively assess whether the LLM use case satisfies the counterfactual invariance property defined in 2.2. One such example of $\upsilon$ is exact match (Rajpurkar et al., 2016), but (Gallegos et al., 2023) argue that this metric is too strict. This work introduces three, less stringent, counterfactual similarity metrics: Counterfactual ROUGE-L, Counterfactual BLEU, and Counterfactual Cosine Similarity, which are extensions of state-of-the-art text similarity metrics (Minaee et al., 2024; Lin, 2004; Papineni et al., 2002; Singhal and Google, 2001; Gomaa and Fahmy, 2013). While the first two assess similarity using token-sequence overlap, the the third assesses similarity using sentence embeddings. All three metrics range in values from 0 to 1, with larger values indicating a greater degree of fairness.
@@ 3.1.3.2 Counterfactual Sentiment Bias
$\mathcal{W}_{1}=\mathbb{E}_{\tau\sim\mathcal{U}(0,1)}|P(Sm(\hat{Y}^{\prime})>\tau)-P(Sm(\hat{Y}^{\prime\prime})>\tau)|,$
@@
This study presents WCSP, defined as the discrepancy in predicted sentiment rates by a sentiment classifier when applied to counterfactually generated LLM outputs, given a threshold $\tau$ for binarizing sentiment scores.
$WCSP=\frac{1}{N}\sum_{i=1}^{N}|P(Sm(\hat{Y}_{i}^{\prime})>\tau)-P(Sm(\hat{Y}_{i}^{\prime\prime})>\tau)|.$

### v2

@@ 1 Introduction
The versatility of current Large Language Models (LLMs) in handling various tasks (Minaee et al., 2024; Liu et al., 2023; Ray, 2023) presents challenges when it comes to evaluating bias and fairness at the model level (Anthis et al., 2024). Existing approaches primarily focus on assessing risk using benchmark data sets containing predefined prompts (Gehman et al., 2020; Dhamala et al., 2021; Nozza et al., 2021; Smith et al., 2022; Parrish et al., 2021; Li et al., 2020; Wang et al., 2024a), masked tokens (Zhao et al., 2018; Rudinger et al., 2018; Nadeem et al., 2021; Levy et al., 2021), or unmasked sentences (Nangia et al., 2020; Barikeri et al., 2021; Jiao et al., 2023; Felkner et al., 2023), assuming that these adequately capture specific bias or fairness risks (Gallegos et al., 2023). However, these assessments are likely to overestimate the risk for use cases where the population of prompts is low risk (Wang et al., 2024b). Moreover, to the best of the author’s knowledge, the current literature does not provide a framework for effectively aligning LLM use cases with suitable metrics for evaluating bias and fairness.
@@
This work aims to address these limitations by developing an actionable LLM bias and fairness evaluation framework defined at the use case level. Drawing inspiration from the classification fairness framework proposed by Saleiro et al. (2018), the framework proposed in this work enables practitioners to map an LLM use case to an appropriate set of bias and fairness evaluation metrics by considering relevant characteristics of the use case and stakeholder values. This evaluation approach is unique in that it incorporates actual prompts from the practitioner’s use case, taking into account the prompt-specific risks that have been demonstrated to significantly increase the likelihood of biased and unfair outcomes (Wang et al., 2024b). By constraining the scope to focused use cases, where prompts are derived from a known population and the task is well-defined, this framework is specifically designed to customize the risk assessment for a specific application.
@@
To introduce the framework, this study first provides formal definitions of bias and fairness desiderata for LLMs from the literature and segments these definitions by risk category. Subsequently, those risks are mapped to a taxonomy of use cases focused on large-scale applications, where human-in-the-loop may be infeasible. Lastly, for each risk category, various bias and fairness evaluation metrics are detailed, with discussions provided on their input requirements, calculation methods, the risks they assess, and circumstances under which they should be applied. As part of this work, a variety of novel bias and fairness metrics are introduced. This includes innovative counterfactual adaptations of recall-oriented understudy for gisting evaluation (ROUGE) (Lin, 2004), bilingual evaluation understudy (BLEU) (Papineni et al., 2002), and cosine similarity (Singhal and Google, 2001), as well a set of stereotype classifier-based metrics that are adapted from analogous toxicity classifier-based metrics.
@@ 2 Bias and Fairness Risks for LLM Use Cases
This section formally defines several pre-requisite terms and concepts upon which subsequent sections rely, several of which are adapted from those provided by Gallegos et al. (2023). [^2] The notation and terminology used in this paper differs from those used by Gallegos et al. (2023). The approach outlined below adopts modified versions of these definitions to better fit the tone and purpose. For further reading on biases in LLMs, the reader may refer to (Gallegos et al., 2023; Blodgett et al., 2020; Kumar et al., 2023; Li et al., 2024; Chu et al., 2024; Ferrara, 2023; Ranaldi et al., 2023; Kotek et al., 2023; Wu and Aji, 2023; Li et al., 2023; Ray, 2023; Nozza et al., 2022; Zhuo et al., 2023). The discussions in this paper explore concepts of bias and fairness in relation to an arbitrary ‘protected attribute’, encompassing examples such as sex, race, age, and sexual orientation, among others. Formal establishment of these concepts is achieved through the use of mathematical notation whenever feasible. The aim is to maintain consistency in notation throughout, even if it deviates from the notation used in the referenced papers.
@@ 2.2.3 Counterfactual Fairness
In many contexts, it is undesirable for an LLM to generate substantially different output as a result of different protected attribute words contained in the input prompts, all else equal (Huang et al., 2020; Nozza et al., 2021; Wang et al., 2024b). Following previous work (Huang et al., 2020; Garg et al., 2019), this concept is hereafter referred to as (lack of) counterfactual fairness. Depending on context and stakeholder values, the practitioner may wish to assess an LLM use case for differences in overall content or sentiment resulting from inclusion of different protected attribute words in a prompt. Below, this section provides a formal definition of the corresponding fairness desiderata known as counterfactual invariance, adapted from Gallegos et al. (2023).
@@ 3 Bias and Fairness Evaluation Metrics
Accounting for prompt-specific risk is important to accurately reflect the risk of a specific use case, as Wang et al. (2024b) find in their evaluations that toxicity probability is 26 to 101 times higher when dealing with toxic prompts compared to non-toxic prompts. Moreover, Goldfarb-Tarrant et al. (2020); Delobelle et al. (2021) have demonstrated that evaluation metrics which take into account the LLM’s task provide a more accurate reflection of the associated risk compared to metrics based on embeddings or token probabilities. Accordingly, the objective of this work is to assess the risks associated with a particular use case, taking into consideration not only the LLM used but also the task at hand and the prompt population. Following the characterization of a use case based on a model and a known population of prompts, as established in Section 2.1, each metric definition presented below is contextualized within an evaluation sample of size $N$ drawn from a known population of prompts $\mathcal{P}_{X}$.
@@ 3.1 Bias Metrics for Text Generation and Summarization Use Cases
While Gallegos et al. (2023) further segment generated-text metrics into three subcategories (namely distribution metrics, classifier metrics, and lexicon metrics), this work segments these metrics according to the risk taxonomy outlined in Section 2.2. In particular, generated-text bias and fairness metrics are segmented into the following categories: toxicity metrics, stereotype metrics, and counterfactual fairness metrics. Toxicity metrics leverage a pre-trained toxicity classifier, such as Perspective API [^12] https://perspectiveapi.com, to assign a toxicity score to an LLM’s output (Chowdhery et al., 2022; Lees et al., 2022; Wang et al., 2024b; Bommasani et al., 2023; Gehman et al., 2020). Stereotype metrics assess the relative co-occurrence of stereotype words with protected attribute words (Bordia and Bowman, 2019; Bommasani et al., 2023) or leverage a pre-trained stereotype classifier to assign a stereotype score to an LLM’s output (Zekun et al., 2023). Lastly, counterfactual fairness metrics assess fairness through a causal lens by exploiting variation in mentions of protected attribute groups in otherwise-identical prompts (Huang et al., 2020). Select metrics are detailed in each category below. [^13] Note that the discussion of generated text bias metrics in this paper is not exhaustive. Rather, the intention is to select an appropriate set of evaluation metrics to address each risk to be assessed. Notable metrics omitted from this framework include HONEST (Nozza et al., 2021), score parity (Sicilia and Alikhani, 2023), Full Gen Bias (Smith et al., 2022), regard score (Sheng et al., 2019), demographic representations (Bommasani et al., 2023), psycholinguistic norms (Dhamala et al., 2021), and gender polarity (Dhamala et al., 2021).
@@ 3.1.1 Toxicity Metrics
Toxicity metrics leverage a pre-trained toxicity classifier to assess general toxicity that is not specific to a particular protected attribute. Given a toxicity classifier $T:\mathcal{Y}\xrightarrow[]{}[0,1]$, this section defines toxicity metrics in the context of an LLM use case $(\mathcal{M},\mathcal{P}_{X})$. Specifically, these metrics are evaluated on a set of $m$ responses per prompt, $\{\hat{Y}_{ij}\}_{j=1}^{m}$, generated by $\mathcal{M}$, from a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$. In practice, variation in responses for the same prompt can be achieved by choosing a non-zero temperature for the LLM or by utilizing various random seeds (Wang et al., 2024b). Below, formal definitions of three toxicity metrics are provided: Expected Maximum Toxicity, Toxicity Probability, and Toxic Fraction. All three metrics range in values from 0 to 1, with smaller values indicating a greater degree of fairness.
@@
Note that the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024b).
@@
As with EMT, the standard choice of $m$ for this metric is $m=25$ (Gehman et al., 2020; Wang et al., 2024b).
@@ 3.1.2.1 Co-occurrence-Based Metrics
Given two protected attribute groups $G^{\prime},G^{\prime\prime}$ with associated sets of protected attribute words $A^{\prime},A^{\prime\prime}$, a set of stereotypical words $W$, a set of stop words $\mathcal{S}$, and an LLM use case $(\mathcal{M},\mathcal{P}_{X})$, the full calculation of COBS is as follows:
$cooccur(w,A|\hat{Y})=\sum_{w_{j},w_{k}\in\hat{Y},w_{j}\neq w_{k}}I(w_{j}=w)\cdot I(w_{k}\in A)\cdot\beta^{dist(w_{j},w_{k})}$
$P(w|A)=\frac{\sum_{i=1}^{N}cooccur(w,A|\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\hat{Y}_{i}}cooccur(\tilde{w},A|\hat{Y}_{i})\cdot I(\tilde{w}\notin\mathcal{S}\cup\mathcal{A})}{\sum_{i=1}^{N}\sum_{a\in A}C(a,\hat{Y}_{i})/\sum_{i=1}^{N}\sum_{\tilde{w}\in\hat{Y}_{i}}C(\tilde{w},\hat{Y}_{i})\cdot I(\tilde{w}\notin\mathcal{S}\cup\mathcal{A})}$
$COBS=\frac{1}{|W|}\sum_{w\in W}\log\frac{P(w|A^{\prime})}{P(w|A^{\prime\prime})},$
where $C(x,\hat{Y}_{i})$ denotes the count of $x$ in $\hat{Y}_{i}$ and $dist(w_{j},w_{k})$ denotes the number of tokens between $w_{j}$ and $w_{k}$. Above, the co-occurrence function $cooccur(w,A|\hat{Y})$ computes a weighted count of words from $A$ that are found within a context window centered around $w$, each time $w$ appears in $\hat{Y}$. [^15] Although Bordia and Bowman (2019) introduce two versions of this metric—one with a fixed-context window and another with an infinite context window—only the version with the infinite context window is incorporated into this framework. In their work, Bordia and Bowman (2019) use $\beta=0.95.$ Note that the functions $cooccur(\tilde{w},A|\hat{Y}_{i})$ and $C(\tilde{w},\hat{Y}_{i})$ are multiplied by zero for $\tilde{w}\in\mathcal{S}\cup\mathcal{A}$ in order to exclude stop words and protected attribute words from these counts. Put simply, COBS computes the relative likelihood that an LLM $\mathcal{M}$ generates output having co-occurrence of $w\in W$ with $A^{\prime}$ versus $A^{\prime\prime}$. This metric has a range of possible values of (,)($\infty$,$\infty$)( - ∞ , ∞ ), with values closer to 0 signifying a greater degree of fairness.
@@
Consider a set of protected attribute groups $\mathcal{G}$, an associated set of protected attribute group lexicons $\mathcal{A}$, and an associated set of stereotypical words $W$. Additionally, let $C(x,\hat{Y})$ denote the number of times that the word $x$ appears in the output $\hat{Y}$, $P^{\text{ref}}$ denote a reference distribution, and $TVD$ denote total variation difference. [^16] The reference distribution recommended by Bommasani et al. (2023) is the uniform distribution. Total variation distance measures the distance between probability distributions. For a given LLM $\mathcal{M}(X;\theta)$ and a sample of prompts $X_{1},...,X_{N}$ drawn from $\mathcal{P}_{X}$, the full computation of SA is as follows:
@@ 3.1.3.1 Counterfactual Similarity
Counterfactual similarity metrics measure the similarity in outputs generated from counterfactual input pairs according to a specified invariance metric $\upsilon$, i.e. $\upsilon(\mathcal{M}(X^{\prime};\theta),\mathcal{M}(X^{\prime\prime};\theta)).$ These metrics effectively assess whether the LLM use case satisfies the counterfactual invariance property defined in Section 2.2. One such example of $\upsilon$ is exact match (Rajpurkar et al., 2016), but (Gallegos et al., 2023) argue that this metric is too strict. This work introduces three, less stringent, counterfactual similarity metrics: Counterfactual ROUGE-L, Counterfactual BLEU, and Counterfactual Cosine Similarity, which are extensions of state-of-the-art text similarity metrics (Minaee et al., 2024; Lin, 2004; Papineni et al., 2002; Singhal and Google, 2001; Gomaa and Fahmy, 2013). While the first two assess similarity using token-sequence overlap, the the third assesses similarity using sentence embeddings. All three metrics range in values from 0 to 1, with larger values indicating a greater degree of fairness.
@@ 3.1.3.2 Counterfactual Sentiment Bias
$SCSP=\mathbb{E}_{\tau\sim\mathcal{U}(0,1)}|P(Sm(\hat{Y}^{\prime})>\tau)-P(Sm(\hat{Y}^{\prime\prime})>\tau)|,$
@@
This study presents WCSP, defined as the difference in predicted sentiment rates by a sentiment classifier applied to counterfactually generated LLM output pairs. Given a threshold $\tau$ for binarizing sentiment scores, this metric is introduced as follows:
$WCSP=\left\lvert\frac{1}{N}\sum_{i=1}^{N}I(Sm(\hat{Y}_{i}^{\prime})>\tau)-\frac{1}{N}\sum_{i=1}^{N}I(Sm(\hat{Y}_{i}^{\prime\prime})>\tau)\right\rvert.$